prettygirl15:

Graph the equation y=4^2+16x-3 1. Direction 2. Y intercept 3. Axis of Symmetry 4. Vertex 5. Zeros

ch3wyc00ki3:

so what direction is it going in first off

prettygirl15:

@ch3wyc00ki3 wrote:
so what direction is it going in first off
i need help on the rest

ch3wyc00ki3:

@prettygirl15 wrote:
@ch3wyc00ki3 wrote:
so what direction is it going in first off
i need help on the rest
ok where does it cross the y axis

Elite:

The equation was graphed correctly by @ch3wyc00ki3. The direction is positive since it goes from left to right. The Y-Intercept (otherwise where the line crosses the Y-axis) is at (0,13). Not enough elaboration in terms of the vertex, you'd need to have shapes on the graph to find those, and here, as far as I can see it's simply a constant line with a positive slope. The zero of a graph (and I assume it to be the real zero here) is basically just the x-intercept, or the opposite of the y-intercept. As such, at least on my graphing calclulator, your real zero should be 0.8125 or (0.8125, 0). That's about as much as I can help you on with the information provided. If there's more to this and I have the entirely wrong idea then tell me, because not many details were provided, you just told us to graph the equation and listed some terms which I assume you want us to find.

surjithayer:

\[y=4x^2+16x-3\] \[y=4(x^2+4x+4-4)-3\]\[=4(x+2)^2-16-3\]\[=4(x+4)^2-19\] 1. it is an upward parabola. 2.y -intercept =-3 3.axis of symmetry is x+4=0 or x=-4 4. vertex : (-4,-19) 5.to find zeros put y=0 \[4x^2+16x-3=0\] \[x=\frac{ -16 \pm \sqrt{16^2-4\times 4\times -3} }{ 2\times4 }\] \[x=\frac{ -16 \pm \sqrt{256+48} }{ 8 }\] \[x=\frac{ -16 \pm \sqrt{304} }{ 8}\] \[x=\frac{ -16 \pm 4\sqrt{19} }{ 8 }\] \[x=-2 \pm\frac{ \sqrt{19}}{2}\]

prettygirl15:

@surjithayer wrote:
\[y=4x^2+16x-3\] \[y=4(x^2+4x+4-4)-3\]\[=4(x+2)^2-16-3\]\[=4(x+4)^2-19\] 1. it is an upward parabola. 2.y -intercept =-3 3.axis of symmetry is x+4=0 or x=-4 4. vertex : (-4,-19) 5.to find zeros put y=0 \[4x^2+16x-3=0\] \[x=\frac{ -16 \pm \sqrt{16^2-4\times 4\times -3} }{ 2\times4 }\] \[x=\frac{ -16 \pm \sqrt{256+48} }{ 8 }\] \[x=\frac{ -16 \pm \sqrt{304} }{ 8}\] \[x=\frac{ -16 \pm 4\sqrt{19} }{ 8 }\] \[x=-2 \pm\frac{ \sqrt{19}}{2}\]
that's for the bottom

27rlyon:

if you dont know the answer just look it up on google its way better than this bruh